# A Simple Model of Current Ramp-Up and Ramp-Down in Tokamaks

### v2

@@ I Introduction
Richard Fitzpatrick’s 2026 article Fitzpatrick:2026 on current ramp-up and ramp-down in tokamaks, confusingly coupled a number of effects that are in principle separable. A disentanglement, Section II, shows his fundamental physics errors in the definition and evolution of the poloidal flux in tokamaks. In particular, Fitzpatrick:
@@
    Did not explicitly include the changes in the flux in the central solenoid, $\psi_{sol}$, in his analysis, which is the only flux that is a directly controllable function of time.
@@
    Assumed a spatially uniform perpendicular energy diffusivity to calculate the current profile that would be expected in a ramp down. Since that particular profile is stable from the 1987 study of TFTR disruptions by Cheng, Furth, and Boozer MHD stab , Fitzpatrick made the remarkable extrapolation that all ramp downs of similar speed are stable.
@@
    Gave a general criterion for avoiding disruptive states, which is clearly incorrect.
Fitzpatrick would have avoided most of these errors had he considered Figure 1 in my article Boozer:flux , and the relation between the poloidal flux enclosed by the magnetic axis of a tokamak and the loop voltage around that axis. As explained just below the figure, this relation follows from the application of Stokes’ Theorem to Faraday’s Law.
@@
Fitzpatrick has insisted on misinterpreting my article as primarily about acceptable ramp-up and ramp-down times. The abstract of his paper states: “there is no indication that the design ramp times are infeasible, as was recently suggested in Boozer (2025 arXiv:2507.05456v1).” Somehow he came to this conclusion when even the first arXiv version, which he emphasizes rather than the version submitted to the Physics of Plasmas, stated the long decay time of the plasma current at fusion temperatures means: “A rapid cooling of the central part of the plasma seems required, but unless this is done with care, the current profile will evolve into a disruptive state.” The same arXiv version also said the restart period in a pulsed tokamak has significant freedom: “In principle, the restart period can be made arbitrarily short by clever choices of the time dependence of $\psi_{sol}(t)$ and the heating power. Of course, the periods in which substantial heating power is required must be short compared to the periods of fusion burn to have net energy to sell.” Fitzpatrick made no note of the sentence in the abstract of the submitted version of my article: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
As discussed in Section II.4, the smallness of the required change in the poloidal flux to cross the full range of stable plasma current profiles implies the current profile in the plasma is determined by the spatial constancy of the loop voltage. Actually Fitzpatrick assumed a constant loop voltage to obtain the single profile of plasma current that he discussed in his Nuclear Fusion paper, but did not explain why this was an accurate approximation.
Despite a number of emails that attempted to correct Fitzpatrick’s mistaken interpretation of my article, he not only persisted in that misinterpretation but reinforced his misinterpretation in every reference to my arXiv article, except in the abstract, by citing “private communication” as well. The only private communication that we had was the emails that tried to correct his misinterpretation.
@@
The reason for Fitzpatrick’s persistence on misinterpreting my article can be explained by his statement of dogma in his September 10, 2025 email, part of his “private communication” reference: “a tokamak plasma is quiescent. There are no large amplitude tearing modes, and the equilibrium profiles are determined by transport rather than MHD activity. The only exception is the sawtooth oscillation, which periodically generates a short burst of MHD activity that prevents the central $q$ value from falling significantly below unity. Given that $q(0)$ is pinned to a value close to unity, and $q(a)$ is fixed by the plasma current, the average broadness of the current profile is easily controlled. Experimentalists have also become adept at further tailoring the current profile by means of off-axis heating. It is not true that there is an inevitable drift to a disruptive state in a tokamak plasma. In fact, tokamak discharges could be maintained indefinitely were it not for the limit in the flux-swing of the central solenoid, and the gradual build up of impurities.”
Fitzpatrick’s dogma seems to imply that disruptions could be easily avoided by a competent machine operator. It was not so simple in JET shutdowns Sozzi:2020 .
Although I have never said that tokamaks have “an inevitable drift to a disruptive state,” ensuring tokamaks have a sufficiently low disruption rate for the economic feasibility of power plants is non-trivial and requires improved understanding.
A similar dogma to the one that prevented Fitzpatrick from acknowledging my views seems to affect the judgement of a significant part of the tokamak community and allows them to overlook even obvious mistakes in science. To correct this, the scientific evidence that does not conform with Fitzpatrick’s statement of dogma needs to be given.
A Physics of Plasmas reviewer of my article said Fitzpatrick’s paper is “a more rigorous model.” A leader in the tokamak program stated that Fitzpatrick’s paper: “refutes your arguments that this (disruptions) inherently dooms tokamaks. This reason goes well beyond Richard’s paper, which attacks the problem from a fundamental and transparent way in a simple model. This is because some of the most sophisticated and benchmarked numerical simulations we have for tokamaks are time-dependent stability/equilibrium codes that tackle the inductive problem. These have been developed over decades of development and testing against thousands of tokamaks discharges on an extremely large range of device sizes, plasma temperatures and engineering design.” Similar comments have come from a number of persons.
It is surprising that anyone believes the quality of the science underlying the design of tokamak power plants is at such a low level that Fitzpatrick’s paper adds important information about disruption avoidance. The glowing descriptions of “sophisticated and benchmarked numerical simulations we have for tokamaks” is an appropriate statement to make to a potential private funder but is not a scientific response to a scientific study into the requirements for avoiding disruptions.
@@
In addition to theory and simulations, empirical studies advance understanding of disruptions and their avoidance. Studies based on the DCAF code developed by Steve Sabbagh’s group have been used to determine Sabbagh:2024 the empirical dependence of tokamak disruptions on plasma quantities such as the internal inductance $\ell_{i}$ and edge safety factor $q_{95}$, Figure 2, The results for MAST-U, which has highly non-circular magnetic surfaces, give preliminary evidence that an internal inductance that is either too low or too high gives disruptive plasmas with a stable region between. An empirical study that is related to that of Figure 2 is given in Figure 3, which is a study of disruptions in JET with an ITER-like wall, JET-dis:2020 . That figure shows that internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ are not sufficient to define a stable region. Are the “sophisticated and benchmarked numerical simulations we have for tokamaks” adequate for explaining Figures 2 and 3?
@@
Ensuring disruptions are sufficiently rare for the feasibility of tokamak power plants would seem to require identifying the externally controllable parameters that can be localized in a way that makes the plasma safe against disruptions. This requirement is more stringent than in existing tokamaks or in JET because of the limitations on diagnostics and actuators in power plants. Generative AI using data from both experiments and simulations could help in this identification.
A major determinant of disruptivity is the profile of the net plasma current $I(\psi_{t})$ enclosed by a magnetic surface that also encloses a toroidal magnetic flux $\psi_{t}$. The derivative $I^{\prime}(\psi_{t})\equiv dI/d\psi_{t})$ is given by the divergence-free part of $\vec{j}_{||}=(j_{||}/B)\vec{B}$, which is the net plasma current, and is determined by the spatial constancy of the loop voltage. Despite the dogma expressed by Fitzpatrick and apparently accepted by many in the tokamak community, the naturally arising profiles of $I(\psi_{t})$ during startup, flattop, and shutdown have many uncertainties. Figures 2 and 3 show that actual tokamak plasmas have many values of $\ell_{i}$, which quantifies the peakedness of the current profile, for a given edge safety factor. There is not just the one current profile that Fitzpatrick focused on.
What theory can do and what I am doing is identifying controllable features in $I(\psi_{t})$ that determine the internal inductance $\ell_{i}$. When the central safety factor $q_{0}$ is held to unity by sawteeth, the derivative $I^{\prime}(0)$ is defined. The determination of the current density $I^{\prime}(\psi_{t})$ by the loop voltage implies $I^{\prime}(\psi_{t})\geq 0$ without negative external current drive. Both a positive $I^{\prime}$ near the edge, as in a H-mode plasma, and triangular shaping of the magnetic surfaces reduce $\ell_{i}$ for a fixed edge to central $q$-factor ratio with an otherwise fixed $I(\psi_{t})$. It would be interesting to determine if extending the empirical data used to construct Figures 2 and 3 would be much more determinate using a four than in a two parameter fit.
The acceptance of a Fitzpatrick-like dogma apparently led to strong negative reactions to two statements that were in the various versions of my arXiv article: The first statement was: “Diversity of concepts is important for the development and optimization of fusion. Nevertheless, an intelligent allocation of resources among the various concepts requires theoretical and computational assessments of the risks, time, and costs of fusion development based on these concepts.” In principle, it would seem anyone who believes in the importance of the rapid development of fusion would agree. The second statement was: “The purpose of this paper is not to remove tokamaks from consideration but rather to determine on what issues research should be focused for tokamaks to be a credible alternative to stellarators. For this, the obvious strategy is to determine what actuators are available to control the things that must be controlled and what plasma data could be available to guide these actuators.” This statement seems controversial only when one believes the external-current determination of stellarator plasmas is not fundamentally different in disruption and other plasma control issues when compared to the effects of internal plasma currents in tokamaks.
Anyone responsible for managing the development of fusion power plants would do well to ponder Admiral Rickover’s well-known quotes based on his success in the launching the first nuclear submarine about fifteen years after the splitting of the uranium nucleus in Berlin. An important quote is: “It is a human inclination to hope things will work out, despite evidence or doubt to the contrary. A successful manager must resist this temptation.”
@@ II.1 Basic properties the poloidal flux
Although ignored by Fitzpatrick, fundamental properties of the poloidal flux and its evolution were illustrated in Figure 1, from my article Boozer:flux , and a simple application of Stokes’ Theorem to Faraday’s Law at the circular magnetic axis of tokamak was given just below that figure.
@@
The poloidal flux $\psi_{p}^{f^{\prime}}$ is related to Fitzpatrick’s definition in which he used a cylindrical $(r,\theta,z)$ model for the poloidal magnetic field in a tokamak.
Fitzpatrick’s definition of poloidal flux differed in two ways from $\psi_{p}^{f^{\prime}}$ of Equation 9. First, he included a factor of no fundamental importance due to the non-circularity of the magnetic surfaces, which for simplicity will be ignored here. The effect of magnetic-surface shape on the rotational transform is given by Equation (237) in Reference NF review . Second and what is of fundamental importance, Fitzpatrick did not include the poloidal flux that lies outside of the plasma but is produced by the plasma current; he retained only the poloidal flux internal to the plasma,
$\psi_{p}^{in}=2\pi R_{ax}\int_{0}^{a}B_{\theta}dr.$ (10)
He did include a time dependent value within his definition of the poloidal flux at the magnetic axis, Equation (14).
For the circular surfaces of a $(r,\theta,z)$ model, $B_{\theta}=\mu_{0}I(r)/(2\pi r)$, where $I(r)=2\pi\int_{0}^{r}j_{t}(r)rdr$ and $j_{t}$ is the net toroidal current density. Fitzpatrick set the contribution to $\psi_{p}^{f^{\prime}}$ for $r>a$ equal to zero, where $a$ is the plasma radius. In a cylindrical approximation, the poloidal flux that lies between $a$ and an outer radius $b>a$ is $\{\mu_{0}I_{p}/(2\pi)\}\int_{a}^{b}dr/r$, which is proportional to $\ln(b/a)$ and becomes logarithmically infinite for $b>>a$.
As noted in my article and stated in many other places, the poloidal flux of a large aspect ratio toroidal shell with major and minor radii $R$ and $a$ carrying a current $I_{p}$ is
@@
For standard tokamaks, the external part of the poloidal flux produced by the plasma current is larger than the flux given by the plasma current that lies within the plasma, the $\psi_{p}^{in}$ of Equation (10). In his Equation (21), Fitzpatrick gives the typical poloidal flux by his calculation as $\psi_{0}=2\pi R_{ax}aB_{\theta}(a)$. When one ignores the increased current required for a given flux due to non-circularity, which affects both $\psi_{p}^{in}$ and $\psi_{p}^{ex}$. Fitzpatrick’s typical flux is $\psi_{0}=\mu_{0}R_{ax}I_{p}$, and the ratio $\psi_{p}^{ex}/\psi_{0}=\ln(8R/a)-2\approx 1.178$ when $R/a=3$. For a parabolic current profile, the ratio of the external to the internal flux produced by the plasma current is even larger, $\psi_{p}^{ex}/\psi_{p}^{in}=1.57$.
@@
II.2 Disentangling Fitzpatrick’s expressions
Richard Fitzpatrick entangles his arguments about poloidal flux evolution with issues that are in principle separable, such as heat transport. This entanglement produces confusion—apparently to Fitzpatrick himself—and a disentanglement is required.
The first step in a disentanglement is determining what is precisely Fitzpatrick’s definition of the poloidal flux and what basic equations it obeys in the cylindrical $(r,\theta,z)$ approximation that he utilizes. His Equations (6) and (9) together with his Equation (37) applied at $\hat{r}\equiv r/a=0$, where $r=a$ is the plasma edge, imply that the Fitzpatrick definition of the poloidal flux $\psi_{f}(r,t)$ is

### v3

@@ I Introduction
Richard Fitzpatrick’s 2026 Nuclear Fusion paper Fitzpatrick:2026 on current ramp-up and ramp-down in tokamaks, confusingly coupled a number of effects that are in principle separable. A disentanglement, Section II, shows the paper contains fundamental physics errors in the definition and evolution of the poloidal flux in tokamaks. In particular, that paper Fitzpatrick:2026 :
@@
    Did not explicitly include the changes in the flux in the central solenoid, $\psi_{sol}$, which is the only flux that is a directly controllable function of time.
@@
    Assumed a time-independent profile profile, which was a spatial constant, of the thermal energy diffusivity to calculate the current profile that would be expected in a ramp down. Since that particular profile is stable from the 1987 study of TFTR disruptions by Cheng, Furth, and Boozer MHD stab , a remarkable extrapolation was made that all ramp downs of similar speed are stable.
@@
    Assumed that the profile of the poloidal magnetic field was essentially independent of time. Constancy would imply that if the initial current profile were stable to disruptions, it would remain stable.
The Nuclear Fusion paper Fitzpatrick:2026 was inspired by an arXiv article by Allen Boozer Boozer:flux and would have avoided most of these errors had Figure 1, which was in the arXiv article, been considered. This figure illustrates the relation between the poloidal flux enclosed by the magnetic axis of a tokamak and the loop voltage around that axis. As explained in the arXiv article Boozer:flux just below the figure, this relation follows from the application of Stokes’ Theorem to Faraday’s Law.
@@
The abstract to the Nuclear Fusion paper Fitzpatrick:2026 states: “there is no indication that the design ramp times are infeasible, as was recently suggested in Boozer (2025 arXiv:2507.05456v1).” Even the first version of the arXiv article Boozer:flux , which is the focus of the paper Fitzpatrick:2026 , stated the long decay time of the plasma current at fusion temperatures means: “A rapid cooling of the central part of the plasma seems required, but unless this is done with care, the current profile will evolve into a disruptive state.” The same arXiv version also said the restart period in a pulsed tokamak has significant freedom: “In principle, the restart period can be made arbitrarily short by clever choices of the time dependence of $\psi_{sol}(t)$ and the heating power. Of course, the periods in which substantial heating power is required must be short compared to the periods of fusion burn to have net energy to sell.” The Nuclear Fusion paper Fitzpatrick:2026 makes no note of the sentence in the abstract of the submitted September 9, 2025 version of the arXiv article Boozer:flux : “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
As discussed in Section II.4, the smallness of the required change in the poloidal flux to cross the full range of stable plasma current profiles implies that current profiles in the plasma, when stable, are determined by the spatial constancy of the loop voltage. Just before Equation (42) of Fitzpatrick:2026 , an even stronger assumption was made. It was assumed that the profile of the poloidal field remains essentially fixed in time, which implies a spatially constant loop voltage. A constant poloidal-field profile ensures that if the original current profile is disruption stable, then it will remain stable. However, a spatially constant loop voltage does not imply the profile of the poloidal field remains essentially fixed in time since the resistivity, on which the loop voltage depends, is a function of the electron temperature $T_{e}$ and the $Z_{eff}$ due to impurities. The profiles of both $T_{e}$ and $Z_{eff}$ are generally time dependent.
@@
Despite a number of emails that attempted to correct the mistaken interpretation of the arXiv article Boozer:flux , before the Nuclear Fusion paper Fitzpatrick:2026 was published, the published version reinforced that misinterpretation by the citations “private communication” in every reference to the arXiv article Boozer:flux , except in the abstract. Each “private communication” was an email. An insight into the view expressed in the Nuclear Fusion paper Fitzpatrick:2026 can be gained from the “private communication” in a September 10, 2025 email: “a tokamak plasma is quiescent. There are no large amplitude tearing modes, and the equilibrium profiles are determined by transport rather than MHD activity. The only exception is the sawtooth oscillation, which periodically generates a short burst of MHD activity that prevents the central $q$ value from falling significantly below unity. Given that $q(0)$ is pinned to a value close to unity, and $q(a)$ is fixed by the plasma current, the average broadness of the current profile is easily controlled. Experimentalists have also become adept at further tailoring the current profile by means of off-axis heating. It is not true that there is an inevitable drift to a disruptive state in a tokamak plasma. In fact, tokamak discharges could be maintained indefinitely were it not for the limit in the flux-swing of the central solenoid, and the gradual build up of impurities.” This reads as if disruptions can be easily avoided by a competent machine operator, but this was not the case in JET shutdowns Sozzi:2020 .
Although the arXiv article Boozer:flux did not imply that tokamaks have “an inevitable drift to a disruptive state,” ensuring tokamaks have a sufficiently low disruption rate for the economic feasibility of power plants is non-trivial and requires improved understanding. Ensuring disruptions are sufficiently rare for the feasibility of tokamak power plants would seem to require the identification of the externally controllable plasma parameters that ensure the plasma is safe against disruptions. This requirement is more stringent in power plants than in existing tokamaks or in JET because of the limitations on diagnostics and actuators. Generative AI using data from both experiments and simulations could help in this identification.
@@
Empirical studies are essential for identifying parameter regions of low disruptivity. Studies based on the DCAF code developed by Steve Sabbagh’s group have been used to determine Sabbagh:2024 the empirical dependence of tokamak disruptions on plasma quantities such as the internal inductance $\ell_{i}$ and edge safety factor $q_{95}$, Figure 2, The results for MAST-U, which has highly non-circular magnetic surfaces, gave preliminary evidence that an internal inductance that is either too low or too high gives disruptive plasmas with a stable region between. An empirical study that is related to that of Figure 2 is given in Figure 3, which is a study of disruptions in JET with an ITER-like wall, JET-dis:2020 . That figure shows that internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ are not sufficient to define a stable region. Are the numerical simulations of tokamaks sufficiently sophisticated and reliable to explain Figures 2 and 3?
@@
A major determinant of disruptivity is the profile of the net plasma current $I(\psi_{t})$ enclosed by a magnetic surface that also encloses a toroidal magnetic flux $\psi_{t}$. The derivative $I^{\prime}(\psi_{t})\equiv dI/d\psi_{t})$ is given by the divergence-free part of $\vec{j}_{||}=(j_{||}/B)\vec{B}$, which is the net plasma current, and is determined by the spatial constancy of the loop voltage. Despite the view expressed in the September 10, 2025 “private communication” quoted above, the naturally arising profiles of $I(\psi_{t})$ during startup, flattop, and shutdown have many uncertainties. Figures 2 and 3 show that actual tokamak plasmas have many values of $\ell_{i}$, which quantifies the peakedness of the current profile, for a given edge safety factor. There is not just the one current profile that the Nuclear Fusion paper Fitzpatrick:2026 focussed on.
What theory can do is identify controllable features in $I(\psi_{t})$ that determine the internal inductance $\ell_{i}$. When the central safety factor $q_{0}$ is held to unity by sawteeth, the derivative $I^{\prime}(0)$ is defined within a factor determined by the ellipticity of the near-axis magnetic surfaces. The determination of the current density $I^{\prime}(\psi_{t})$ by a spatially-constant loop voltage implies $I^{\prime}(\psi_{t})\geq 0$ without negative external current drive. Both a positive $I^{\prime}$ near the edge, as in a H-mode plasma, and triangular shaping of the magnetic surfaces reduce $\ell_{i}$ for a fixed edge to central $q$-factor ratio with an otherwise fixed $I(\psi_{t})$. It is important to determine if extending the empirical data used to construct Figures 2 and 3 from a two parameter to a four parameter fit would define regions that have a sufficiently robust stability against disruptions.
@@ II.1 Basic properties the poloidal flux
Although ignored in the Nuclear Fusion paper Fitzpatrick:2026 , fundamental properties of the poloidal flux and its evolution were illustrated in Figure 1, from the arXiv article Boozer:flux . Just below that figure, the properties of the evolution of the poloidal flux enclosed by the circular magnetic axis of a tokamak were obtained by a simple application of Stokes’ Theorem to Faraday’s Law.
@@
The poloidal flux $\psi_{p}^{f^{\prime}}$ is related to the definition in the paper Fitzpatrick:2026 , which used a cylindrical $(r,\theta,z)$ model for the poloidal magnetic field in a tokamak.
The definition of poloidal flux in Fitzpatrick:2026 differed in two ways from $\psi_{p}^{f^{\prime}}$ of Equation 9. First, a factor of no fundamental importance due to the non-circularity of the magnetic surfaces was included, which for simplicity will be ignored. The effect of magnetic-surface shape on the rotational transform is given by Equation (237) in Reference NF review . Second and what is of fundamental importance, the paper Fitzpatrick:2026 did not include the poloidal flux that lies outside of the plasma but is produced by the plasma current; only the poloidal flux internal to the plasma was retained,
$\psi_{p}^{in}=2\pi R_{ax}\int_{0}^{a}B_{\theta}dr,$ (10)
where $a$ is the plasma radius. An additive time dependent value of the poloidal flux at the magnetic axis, $\psi_{p}^{ax}(t)$, was included in the poloidal flux evolution, see Equation (14), but the time dependence of the poloidal flux at the axis, unlike the poloidal flux in the central solenoid is not directly controlable.
For the circular surfaces of a $(r,\theta,z)$ model, $B_{\theta}=\mu_{0}I(r)/(2\pi r)$, where $I(r)=2\pi\int_{0}^{r}j_{t}(r)rdr$ and $j_{t}$ is the net toroidal current density. The paper Fitzpatrick:2026 set the plasma contribution to the poloidal flux equal to zero for $r>a$. In a cylindrical approximation, the poloidal flux that lies between the radius of the plasma $a$ and an outer radius $b>a$ is $\{\mu_{0}I_{p}/(2\pi)\}\int_{a}^{b}dr/r$, which is proportional to $\ln(b/a)$ and becomes logarithmically infinite for $b>>a$.
As noted in the arXiv article Boozer:flux and stated in many other places, the poloidal flux of a large aspect ratio toroidal shell with major and minor radii $R$ and $a$ carrying a current $I_{p}$ is
@@
For standard tokamaks, the external part of the poloidal flux produced by the plasma current is larger than the flux given by the plasma current that lies within the plasma, the part $\psi_{p}^{in}$ of Equation (10). Equation (21) of the Nuclear Fusion paper Fitzpatrick:2026 states the typical poloidal flux is $\psi_{0}=2\pi R_{ax}aB_{\theta}(a)$. When one ignores the increased current required for a given flux due to non-circularity, which affects both $\psi_{p}^{in}$ and $\psi_{p}^{ex}$, Fitzpatrick’s typical flux is $\psi_{0}=\mu_{0}R_{ax}I_{p}$. The ratio $\psi_{p}^{ex}/\psi_{0}=\ln(8R/a)-2\approx 1.178$ when $R/a=3$. For a parabolic current profile, the ratio of the external to the internal flux produced by the plasma current is even larger, $\psi_{p}^{ex}/\psi_{p}^{in}=1.57$.
@@
II.2 Disentangling the expressions used in Fitzpatrick:2026
The paper Fitzpatrick:2026 entangles arguments about poloidal flux evolution with issues that are in principle separable, such as heat transport. This entanglement produces confusion—apparently in Fitzpatrick:2026 itself—and a disentanglement is required.
The first step in a disentanglement is determining what is the precise definition of the poloidal flux in Fitzpatrick:2026 and what are the equations it obeys in the cylindrical $(r,\theta,z)$ approximation that is utilized. Equations (6) and (9) in Fitzpatrick:2026 together with Equation (37) applied at $\hat{r}\equiv r/a=0$, where $r=a$ is the plasma edge, imply that the definition of the poloidal flux $\psi_{f}(r,t)$ is